# Asymptotic Distribution-free Change-point Detection for Modern Data Based on a New Ranking Scheme

### v3

@@ III-A Asymptotic null distributions of the basic processes
$\big{\{}Z_{\mathrm{diff}}(\lfloor nu\rfloor):0<u<1\big{\}}\text{ and }\big{\{}Z_{w}(\lfloor nu\rfloor):0<u<1\big{\}}$ (9)
for the single change-point alternative and
$\big{\{}Z_{\mathrm{diff}}(\lfloor nu\rfloor,\lfloor nv\rfloor):0<u<v<1\big{\}}\text{ and }\big{\{}Z_{w}(\lfloor nu\rfloor,\lfloor nv\rfloor):0<u<v<1\big{\}}$ (10)
@@
Conditions (1)-(6) put restrictions on the similarity graph and the ranks. Importantly, they require that there should not be an excessive number of hub nodes in the graph and that the variation of the average row-wise ranks, $V_{r}$, should not be dominated by a small portion of elements. Specifically, when the largest degree of $G_{k}$ is bounded by $Ck$ for some constant $C$, Conditions (1), (2), (4), and (6) always hold for $k$-NNG and $k$-MST, and Conditions (3) and (5) also tend to be satisfied as verified by simulation studies. More detailed discussions can be found in Appendix E of [48]. Note that these conditions are the same as [48], which is exciting because we do not need extra conditions when we extend the statistics from the two-sample setting to the change-point detection setting.
@@ III-B Tail probabilities
Based on Theorems 1 and 2, we are able to approximate the tail probabilities by utilizing Woodroofe’s method [54, 55] and Siegmund’s method [56, 57]. Specifically, following the routine of [30], and the proof of Proposition 3.4 in [28], we can approximate the tail probabilities by
${\mathbbm{P}}\big{(}\max_{n_{0}\leq t\leq n_{1}}M_{R}(t)>b\big{)}\approx 1-{\mathbbm{P}}\big{(}\max_{n_{0}\leq t\leq n_{1}}Z_{w}(t)<b\big{)}{\mathbbm{P}}\big{(}\max_{n_{0}\leq t\leq n_{1}}|Z_{\mathrm{diff}}(t)|<b\big{)}\,,$ (11)
${\mathbbm{P}}\big{(}\max_{1\leq t_{1}<t_{2}\leq n\atop l_{0}\leq t_{2}-t_{1}\leq l_{1}}M_{R}(t_{1},t_{2})>b\big{)}\approx 1-{\mathbbm{P}}\big{(}\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{w}(t_{1},t_{2})<b\big{)}{\mathbbm{P}}\big{(}\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}|Z_{\mathrm{diff}}(t_{1},t_{2})|<b\big{)}\,,$ (12)
@@
${\mathbbm{P}}\big{(}\max_{n_{0}\leq t\leq n_{1}}Z_{w}(t)>b\big{)}\approx b\phi(b)\int_{\frac{n_{0}}{n}}^{\frac{n_{1}}{n}}h_{w}(n,x)\nu\big{(}b\sqrt{2h_{w}(n,x)/n}\big{)}{\rm d}x\,,$ (13)
${\mathbbm{P}}\big{(}\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{w}(t_{1},t_{2})>b\big{)}\approx b^{3}\phi(b)\int_{\frac{l_{0}}{n}}^{\frac{l_{1}}{n}}\Big{(}h_{w}(n,x)\nu\big{(}b\sqrt{2h_{w}(n,x)/n}\big{)}\Big{)}^{2}(1-x){\rm d}x\,,$ (14)
${\mathbbm{P}}\big{(}\max_{n_{0}\leq t\leq n_{1}}Z_{\mathrm{diff}}(t)>b\big{)}\approx b\phi(b)\int_{\frac{n_{0}}{n}}^{\frac{n_{1}}{n}}h_{\mathrm{diff}}(n,x)\nu\big{(}b\sqrt{2h_{\mathrm{diff}}(n,x)/n}\big{)}{\rm d}x\,,$ (15)
${\mathbbm{P}}\big{(}\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{\mathrm{diff}}(t_{1},t_{2})>b\big{)}\approx b^{3}\phi(b)\int_{\frac{l_{0}}{n}}^{\frac{l_{1}}{n}}\Big{(}h_{\mathrm{diff}}(n,x)\nu\big{(}b\sqrt{2h_{\mathrm{diff}}(n,x)/n}\big{)}\Big{)}^{2}(1-x){\rm d}x\,.$ (16)
@@
$h_{w}(n,x)=\frac{(n-1)(2nx^{2}-2nx+1)}{2x(1-x)(nx-1)(nx-n+1)}\quad\text{ and }\quad h_{\mathrm{diff}}(n,x)=\frac{1}{2x(1-x)}\,.$
Here $v(x)$ is approximated as $v(x)\approx(2/x)(\Phi(x/2)-0.5)/\big{(}(x/2)\Phi(x/2)+\phi(x/2)\big{)}$ [58], where $\Phi(\cdot)$ and $\phi(\cdot)$ are the standard normal cumulative density function and standard normal density function, respectively.
@@ III-C Skewness correction
As observed by [28, 30], the analytical approximations of (11) and (12) can be improved by skewness correction when $n_{0}$ and $n-n_{1}$ decrease. It can be seen clearly in Figure 3 that $Z_{w}(t)$ and $Z_{\mathrm{diff}}(t)$ are more skewed toward the two ends. To be specific, instead of using (13)-(16) to approximate (11) and (12), we use
${\mathbbm{P}}\big{(}\max_{n_{0}\leq t\leq n_{1}}Z_{w}(t)>b\big{)}\approx b\phi(b)\int_{\frac{n_{0}}{n}}^{\frac{n_{1}}{n}}K_{w}(nx)h_{w}(n,x)\nu\big{(}b\sqrt{2h_{w}(n,x)/n}\big{)}{\rm d}x\,,$ (17)
${\mathbbm{P}}\big{(}\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{w}(t_{1},t_{2})>b\big{)}\approx b^{3}\phi(b)\int_{\frac{l_{0}}{n}}^{\frac{l_{1}}{n}}K_{w}(nx)\Big{(}h_{w}(n,x)\nu\big{(}b\sqrt{2h_{w}(n,x)/n}\big{)}\Big{)}^{2}(1-x){\rm d}x\,,$ (18)
${\mathbbm{P}}\big{(}\max_{n_{0}\leq t\leq n_{1}}Z_{\mathrm{diff}}(t)>b\big{)}\approx b\phi(b)\int_{\frac{n}{n}}^{\frac{n_{1}}{n}}K_{\mathrm{diff}}(nx)h_{\mathrm{diff}}(n,x)\nu\big{(}b\sqrt{2h_{\mathrm{diff}}(n,x)/n}\big{)}{\rm d}x\,,$ (19)
${\mathbbm{P}}\big{(}\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{\mathrm{diff}}(t_{1},t_{2})>b\big{)}\approx b^{3}\phi(b)\int_{\frac{l_{0}}{n}}^{\frac{l_{1}}{n}}K_{\mathrm{diff}}(nx)\Big{(}h_{\mathrm{diff}}(n,x)\nu\big{(}b\sqrt{2h_{\mathrm{diff}}(n,x)/n}\big{)}\Big{)}^{2}(1-x){\rm d}x\,,$ (20)
where for $j=w,\mathrm{diff}$,
$K_{j}(t)=\frac{\exp\Big{(}\frac{1}{2}\big{(}b-\hat{\theta}_{b,j}(t)\big{)}^{2}+\frac{1}{6}\gamma_{j}(t)\hat{\theta}_{b,j}(t)^{3}\Big{)}}{\sqrt{1+\gamma_{j}(t)\hat{\theta}_{b,j}(t)}}$
with $\hat{\theta}_{b,j}(t)=\frac{-1+\sqrt{1+2\gamma_{j}(t)b}}{\gamma_{j}(t)}$ and $\gamma_{j}(t)={\mathbbm{E}}\big{(}Z_{j}^{3}(t)\big{)}$. The only unknown quantities in the above expressions are $\gamma_{w}(t)$ and $\gamma_{\mathrm{diff}}(t)$, whose exact analytic expressions are quite long and are thus provided in Supplement S3.
@@
(a) $y_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N_{100}(\mathbf{0}_{100},\mathbf{I}_{100})$ (b) $y_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}t_{5}(\mathbf{0}_{100},\mathbf{I}_{100})$
@@ III-D Assessment of finite sample approximations
Here we assess the performance of the asymptotic approximations with finite samples. For a constant $a$, we define the first-order auto-regressive correlation matrix $\bm{\Sigma}(a)=(a^{|i-j|})_{i,j=1}^{d}\in\mathbb{R}^{d\times d}$. We consider three distributions for three different dimensions $d=20,100$ and $1000$ with $n=1000$:
@@
    the multivariate Gaussian distribution $y_{i}\sim N_{d}\big{(}\mathbf{0}_{d},\bm{\Sigma}(0.6)\big{)}$;
@@
    the multivariate $t_{5}$ distribution $y_{i}\sim t_{5}\big{(}\mathbf{0}_{d},\bm{\Sigma}(0.5)\big{)}$;
@@
    the multivariate log-normal distribution $y_{i}\sim\exp\big{(}N_{d}\big{(}\mathbf{0}_{d},\bm{\Sigma}(0.4)\big{)}\big{)}$.
We report the empirical sizes estimated by $1,000$ Monte Carlo simulations. Here, we focus on the graph-induced rank in $k$-NNG. We denote the scan statistic $M_{R}(t)$ on the graph-induced rank in $k$-NNG by $\rm M_{g}$-NN. We set $n_{1}=n-n_{0}$ and consider $n_{0}=[0.025n],[0.05n],[0.1n]$. The nominal level is set to be $0.05$. We see that the empirical sizes are well controlled under all settings even for $n_{0}$ as small as $[0.025n]$ (Table I). We further present the empirical sizes of the $\rm M_{g}$-NN in Figure 4, considering $k=[n^{\lambda}]$ with $\lambda$ ranging from $0.2$ to $0.8$ in $k$-NNG. This visualization indicates that the empirical sizes are close to the nominal levels across a broad spectrum of $\lambda$ values.
@@ III-E Consistency
$M(\delta_{1},\delta_{2})=\lim_{n\rightarrow\infty}\frac{M_{R}\big{(}[\delta_{1}n],[\delta_{2}n]\big{)}}{\sqrt{n}}\text{ for }0\leq\delta_{1}<\delta_{2}<1\text{ and }M(\delta)=M(0,\delta)\text{ for }0<\delta<1\,.$
@@
    $\sup_{\delta\in(0,1)}\Big{|}\frac{M_{R}([\delta n])}{\sqrt{n}}-M(\delta)\Big{|}\stackrel{{\scriptstyle P}}{{\rightarrow}}0\text{ as }n\rightarrow\infty\,,$ (21)
    where $\stackrel{{\scriptstyle P}}{{\rightarrow}}$ denotes the convergence in probability. Then the scan statistic of $M_{R}(t)$ is consistent in that it will reject $H_{0}$ against $H_{1}$ with probability going to one for any significance level $10<\alpha<$ 1 and
@@
    $\sup_{0<\delta_{1}<\delta_{2}<1}\Big{|}\frac{M_{R}([\delta_{1}n],[\delta_{2}n])}{\sqrt{n}}-M(\delta_{1},\delta_{2})\Big{|}\stackrel{{\scriptstyle P}}{{\rightarrow}}0\text{ as }n\rightarrow\infty\,,$ (22)
    then the scan statistic of $M_{R}(t_{1},t_{2})$ is consistent in that it will reject $H_{0}$ against $H_{2}$ with probability going to one for any significance level $10<\alpha<$ 1 and
@@
    the multivariate Gaussian distribution $\big{(}N_{d}(\mathbf{0}_{d},\mathbf{I}_{d}),N_{d}(0.1\mathbf{1}_{d},\mathbf{I}_{d})\big{)}$;
@@
    the multivariate $t_{3}$ distribution $\big{(}t_{3}(\mathbf{0}_{d},\mathbf{I}_{d}),t_{3}(0.1\mathbf{1}_{d},1.02^{2}\mathbf{I}_{d}\big{)}$;
@@
    the multivariate Cauchy distribution $\big{(}{\rm Cauchy}_{d}(\mathbf{0}_{d},\mathbf{I}_{d}),{\rm Cauchy}_{d}(2\mathbf{1}_{d},\mathbf{I}_{d})\big{)}$.
@@
Figure 5: Ten independent sequences (represented by different colors) of $M_{R}([\delta n])/\sqrt{n}$ against $\delta$ for $n=200,800,1600$ and $6400$ for the three settings.

### v4

@@ III-A Asymptotic null distributions of the basic processes
$\big\{Z_{\mathrm{diff}}(\lfloor nu\rfloor):0<u<1\big\}\text{ and }$ (10)
$\big\{Z_{w}(\lfloor nu\rfloor):0<u<1\big\}$
for the single change-point alternative, and
$\big\{Z_{\mathrm{diff}}(\lfloor nu\rfloor,\lfloor nv\rfloor):0<u<v<1\big\}\text{ and }$ (11)
$\big\{Z_{w}(\lfloor nu\rfloor,\lfloor nv\rfloor):0<u<v<1\big\}$
@@
Conditions (1)-(6) put restrictions on the similarity graph and the ranks. Importantly, they require that there should not be an excessive number of hub nodes in the graph and that the variation of the average row-wise ranks, $V_{r}$, should not be dominated by a small portion of elements. Specifically, when the largest degree of $G_{k}$ is bounded by $Ck$ for some constant $C$, Conditions (1), (2), (4), and (6) always hold for $k$-NNG and $k$-MST, and Conditions (3) and (5) also tend to be satisfied as verified by simulation studies. More detailed discussions can be found in Appendix E of [rise]. Note that these conditions are the same as [rise], which is exciting because we do not need extra conditions when we extend the statistics from the two-sample setting to the change-point detection setting.
@@ III-B Tail probabilities
Based on Theorems 1 and 2, we are able to approximate the tail probabilities using Woodroofe’s method [woodroofe1976frequentist, woodroofe1978large] and Siegmund’s method [siegmund1988approximate, siegmund1992tail]. Specifically, following the routine of [chu2019asymptotic], and the proof of Proposition 3.4 in [chen2015graph], we can approximate the tail probabilities by
${\mathbbm{P}}\big(\max_{n_{0}\leq t\leq n_{1}}M_{R}(t)>b\big)$ (12)
$\approx 1-{\mathbbm{P}}\big(\max_{n_{0}\leq t\leq n_{1}}Z_{w}(t)<b\big){\mathbbm{P}}\big(\max_{n_{0}\leq t\leq n_{1}}|Z_{\mathrm{diff}}(t)|<b\big),$
$\approx 1-{\mathbbm{P}}\big(\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{w}(t_{1},t_{2})<b\big)$
$\hskip 28.45274pt\times{\mathbbm{P}}\big(\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}|Z_{\mathrm{diff}}(t_{1},t_{2})|<b\big),$
@@
${\mathbbm{P}}\big(\max_{n_{0}\leq t\leq n_{1}}Z_{w}(t)>b\big)$ (14)
$\approx b\phi(b)\int_{\frac{n_{0}}{n}}^{\frac{n_{1}}{n}}h_{w}(n,x)\nu\big(b\sqrt{2h_{w}(n,x)/n}\big){\rm d}x,$
${\mathbbm{P}}\big(\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{w}(t_{1},t_{2})>b\big)$ (15)
$\approx b^{3}\phi(b)\int_{\frac{l_{0}}{n}}^{\frac{l_{1}}{n}}\Big(h_{w}(n,x)\nu\big(b\sqrt{2h_{w}(n,x)/n}\big)\Big)^{2}(1-x){\rm d}x,$
${\mathbbm{P}}\big(\max_{n_{0}\leq t\leq n_{1}}Z_{\mathrm{diff}}(t)>b\big)$ (16)
$\approx b\phi(b)\int_{\frac{n_{0}}{n}}^{\frac{n_{1}}{n}}h_{\mathrm{diff}}(n,x)\nu\big(b\sqrt{2h_{\mathrm{diff}}(n,x)/n}\big){\rm d}x,$
${\mathbbm{P}}\big(\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{\mathrm{diff}}(t_{1},t_{2})>b\big)$ (17)
$\approx b^{3}\phi(b)\int_{\frac{l_{0}}{n}}^{\frac{l_{1}}{n}}\Big(h_{\mathrm{diff}}(n,x)\nu\big(b\sqrt{2h_{\mathrm{diff}}(n,x)/n}\big)\Big)^{2}(1-x){\rm d}x.$
@@
$h_{w}(n,x)$ $=\frac{(n-1)(2nx^{2}-2nx+1)}{2x(1-x)(nx-1)(nx-n+1)},$
$h_{\mathrm{diff}}(n,x)$ $=\frac{1}{2x(1-x)}.$
Here $v(x)$ is approximated as $v(x)\approx\frac{(2/x)(\Phi(x/2)-0.5)}{(x/2)\Phi(x/2)+\phi(x/2)}$ [siegmund2007statistics], where $\Phi(\cdot)$ and $\phi(\cdot)$ are the standard normal cumulative density function and standard normal density function, respectively.
@@ III-C Skewness correction
As observed by [chen2015graph, chu2019asymptotic], the analytical approximations of (12) and (13) can be improved by skewness correction when $n_{0}$ and $n-n_{1}$ decrease. It can be seen clearly in Figure 3 that $Z_{w}(t)$ and $Z_{\mathrm{diff}}(t)$ are more skewed toward the two ends. To be specific, instead of using (14)-(17) to approximate (12) and (13), we use
${\mathbbm{P}}\big(\max_{n_{0}\leq t\leq n_{1}}Z_{w}(t)>b\big)$ (18)
$\approx b\phi(b)\int_{\frac{n_{0}}{n}}^{\frac{n_{1}}{n}}K_{w}(nx)h_{w}(n,x)\nu\big(b\sqrt{2h_{w}(n,x)/n}\big){\rm d}x,$
${\mathbbm{P}}\big(\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{w}(t_{1},t_{2})>b\big)$ (19)
$\approx b^{3}\phi(b)\int_{\frac{l_{0}}{n}}^{\frac{l_{1}}{n}}\Big(h_{w}(n,x)\nu\big(b\sqrt{2h_{w}(n,x)/n}\big)\Big)^{2}$
$\hskip 71.13188pt\times K_{w}(nx)(1-x){\rm d}x,$
${\mathbbm{P}}\big(\max_{n_{0}\leq t\leq n_{1}}Z_{\mathrm{diff}}(t)>b\big)$ (20)
$\approx b\phi(b)\int_{\frac{n}{n}}^{\frac{n_{1}}{n}}K_{\mathrm{diff}}(nx)h_{\mathrm{diff}}(n,x)\nu\big(b\sqrt{2h_{\mathrm{diff}}(n,x)/n}\big){\rm d}x,$
${\mathbbm{P}}\big(\max_{l_{0}\leq t_{2}-t_{1}\leq l_{1}}Z_{\mathrm{diff}}(t_{1},t_{2})>b\big)$ (21)
$\approx b^{3}\phi(b)\int_{\frac{l_{0}}{n}}^{\frac{l_{1}}{n}}\Big(h_{\mathrm{diff}}(n,x)\nu\big(b\sqrt{2h_{\mathrm{diff}}(n,x)/n}\big)\Big)^{2}$
$\hskip 71.13188pt\times K_{\mathrm{diff}}(nx)(1-x){\rm d}x,$
where $K_{j}(t)=\frac{\exp\Big(\frac{1}{2}\big(b-\hat{\theta}_{b,j}(t)\big)^{2}+\frac{1}{6}\gamma_{j}(t)\hat{\theta}_{b,j}(t)^{3}\Big)}{\sqrt{1+\gamma_{j}(t)\hat{\theta}_{b,j}(t)}}$ for $j=w,\mathrm{diff}$, with $\hat{\theta}_{b,j}(t)=\frac{-1+\sqrt{1+2\gamma_{j}(t)b}}{\gamma_{j}(t)}$ and $\gamma_{j}(t)={\mathbbm{E}}\big(Z_{j}^{3}(t)\big)$. The only unknown quantities in the above expressions are $\gamma_{w}(t)$ and $\gamma_{\mathrm{diff}}(t)$, whose exact analytic expressions are quite long and are thus provided in Supplement S3.
@@
(a) $y_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N_{100}(\mathbf{0}_{100},\mathbf{I}_{100})$ (b) $y_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Mt_{5}(\mathbf{0}_{100},\mathbf{I}_{100})$
@@ III-D Assessment of finite sample approximations
Here we assess the performance of the asymptotic approximations with finite samples. For a constant $a$, we define the first-order auto-regressive correlation matrix as $\bm{\Sigma}(a)=(a^{|i-j|})_{i,j=1}^{d}\in\mathbb{R}^{d\times d}$. We consider three distributions for three different dimensions $d=20,100$ and $1000$ with $n=1000$:
@@
    Multivariate Gaussian: $y_{i}\sim N_{d}\big(\mathbf{0}_{d},\bm{\Sigma}(0.6)\big)$;
@@
    Multivariate $t_{5}$: $y_{i}\sim Mt_{5}\big(\mathbf{0}_{d},\bm{\Sigma}(0.5)\big)$;
@@
    Multivariate log-normal: $y_{i}\sim\exp\big(N_{d}\big(\mathbf{0}_{d},\bm{\Sigma}(0.4)\big)\big)$.
We report the empirical sizes estimated by $1,000$ Monte Carlo simulations. Here, we focus on the graph-induced rank in $k$-NNG. We denote the scan statistic $M_{R}(t)$ on the graph-induced rank in $k$-NNG by $\rm M_{g}$-NN. We set $n_{1}=n-n_{0}$ and consider $n_{0}=[0.025n],[0.05n],[0.1n]$. The nominal level is set to be $0.05$. We see that the empirical sizes are well controlled across all settings even for $n_{0}$ as small as $[0.025n]$ (Table I). We further present the empirical sizes of the $\rm M_{g}$-NN in Figure 4, considering $k=[n^{\lambda}]$ with $\lambda$ ranging from $0.2$ to $0.95$ in $k$-NNG. This visualization indicates that the empirical sizes are close to the nominal levels across a broad spectrum of $\lambda$ values.
@@ III-E Consistency
$M(\delta_{1},\delta_{2})=\lim_{n\rightarrow\infty}\frac{M_{R}\big([\delta_{1}n],[\delta_{2}n]\big)}{\sqrt{n}}\text{ for }0\leq\delta_{1}<\delta_{2}<1$
and $M(\delta)=M(0,\delta)$ for $0<\delta<1$.
@@
    $\sup_{\delta\in(0,1)}\Big|\frac{M_{R}([\delta n])}{\sqrt{n}}-M(\delta)\Big|\stackrel{{\scriptstyle P}}{{\rightarrow}}0$ (22)
    as $n\rightarrow\infty$, where $\stackrel{{\scriptstyle P}}{{\rightarrow}}$ denotes the convergence in probability. Then the scan statistic of $M_{R}(t)$ is consistent in that it will reject $H_{0}$ against $H_{1}$ with probability going to one for any significance level $0<\alpha<1$ and
@@
    $\sup_{0<\delta_{1}<\delta_{2}<1}\Big|\frac{M_{R}([\delta_{1}n],[\delta_{2}n])}{\sqrt{n}}-M(\delta_{1},\delta_{2})\Big|\stackrel{{\scriptstyle P}}{{\rightarrow}}0$ (23)
    as $n\rightarrow\infty$, then the scan statistic of $M_{R}(t_{1},t_{2})$ is consistent in that it will reject $H_{0}$ against $H_{2}$ with probability going to one for any significance level $0<\alpha<1$ and
@@
Figure 5: Ten independent sequences (depicted using different colors) of $M_{R}([\delta n])/\sqrt{n}$ against $\delta$ for $n=200,800,1600$ and $6400$ for the three settings.
@@
    Multivariate Gaussian: $\big(N_{d}(\mathbf{0}_{d},\mathbf{I}_{d}),N_{d}(0.1\mathbf{1}_{d},\mathbf{I}_{d})\big)$;
@@
    Multivariate $t_{3}$: $\big(Mt_{3}(\mathbf{0}_{d},\mathbf{I}_{d}),Mt_{3}(0.1\mathbf{1}_{d},1.02^{2}\mathbf{I}_{d}\big)$;
@@
    Multivariate Cauchy: $\big(MC_{d}(\mathbf{0}_{d},\mathbf{I}_{d}),MC_{d}(2\mathbf{1}_{d},\mathbf{I}_{d})\big)$.
@@
Remark 4
The theoretical results in Theorem 3 are derived under the assumption of $k=O(1)$. However, in practice, we have observed that using larger values of $k$, specifically $k=O(n^{\lambda})$ for $\lambda>0$, leads to increased power. This improvement can be attributed to the fact that larger $k$ values incorporate more similarity information into the graph. Although increasing $k$ may introduce additional noise, when $\lambda$ is not too large, the benefits of additional similarity information outweigh the noise, compared to the case where $k=O(1)$. The proof of Theorem 3 builds on the foundational work by [henze1999multivariate], which considered the minimum spanning tree, i.e., $k=1$. If a similar result to that in [henze1999multivariate] can be established for $k=O(n^{\lambda})$ with $\lambda>0$, we believe that the assumption in Theorem 3 could be relaxed to $k=O(n^{\lambda})$ as well. We leave this extension for future research.